OpenStudy (anonymous):

Find all solutions to the equation. sin2x + sin x = 0 @johnweldon1993

Nnesha (nnesha):

take out common factor let sin x=y so y^2 +y what's the common factor ?

OpenStudy (johnweldon1993):

This is \(\large sin^2(x) + sin(x) = 0\) right? or \(\large sin(2x) + sin(x) = 0\) ?

OpenStudy (anonymous):

no first one lol

OpenStudy (johnweldon1993):

Then perfect go ahead with the process @Nnesha was showing :) Look at it as sin(x) = y so sin^2(x) = y^2 \[\large y^2 + y = 0\] What can you factor out?

OpenStudy (xapproachesinfinity):

still troubled by this or solved it

OpenStudy (anonymous):

troubled @xapproachesinfinity

OpenStudy (xapproachesinfinity):

so you have no idea how to factor y^2+y=0? what is common btw y^2 and y?

OpenStudy (xapproachesinfinity):

let's us go to the basics how can you do this a(b+c)=? using distributive law?

OpenStudy (xapproachesinfinity):

i want you to interact with me otherwise i will just take my leave haha

OpenStudy (anonymous):

sorry i was doing somethin away from the laptop.. but i have NO IDEA how to do this ABSOLUTELY NONE!

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

yes?

OpenStudy (xapproachesinfinity):

ok back let's pick it up where i left it know of distributive law?

OpenStudy (anonymous):

nope :( im so sorry im making this hell for u wish i did :/

OpenStudy (xapproachesinfinity):

realize that is the basic math? from lower grades

OpenStudy (xapproachesinfinity):

so distributive law tells us that a(b+c)=ab+ac the left hand side is the same are the right hand side any doubt that it true?

OpenStudy (xapproachesinfinity):

same as*

OpenStudy (xapproachesinfinity):

well?

OpenStudy (anonymous):

wat?

OpenStudy (xapproachesinfinity):

i said do you agree that a(b+c)=ab+ac or against this arithmetic law? the problem here is we need to establish some sense first before we tackle that problem of your that deals with trig

OpenStudy (anonymous):

yes

OpenStudy (xapproachesinfinity):

so you have no problem to tell what this is gonna be bd+be=? using the same law?

OpenStudy (xapproachesinfinity):

oh well! so it is gonna be bd+be=b(d+e) yes?

OpenStudy (anonymous):

yes

OpenStudy (xapproachesinfinity):

now tell me what did i just do?

OpenStudy (xapproachesinfinity):

describe me what i did?

OpenStudy (anonymous):

u factored y^2 and y

OpenStudy (xapproachesinfinity):

no just stay on what i did that would come later! what i'm trying to do here to show you how we can accomplish the factorization for y^2+y

OpenStudy (anonymous):

oh ok got it

OpenStudy (xapproachesinfinity):

i want you to fully understand this! so you will have further problem with such exercice!

OpenStudy (xapproachesinfinity):

so here really what i did is using the that very distributive law of arithmetic to get from bd+be to b(d+e)

OpenStudy (anonymous):

alright got it

OpenStudy (xapproachesinfinity):

ok let me test one more time xy+2x=?

OpenStudy (xapproachesinfinity):

again distributive law!

OpenStudy (anonymous):

nvm

OpenStudy (xapproachesinfinity):

this is all for the sake of you being better at this, if that's the case then it is really fine with me we can gave you the answer, but what is the point of doing it any way

OpenStudy (xapproachesinfinity):

give*

Nnesha (nnesha):

one more try y^2 + y let y = chocolates exponent represents number of chocolates  how many chocolates they both people have n common ?